# A $2$-distance $(2\Delta+7)$-coloring of planar graphs

### v2

@@ 2.2. The case $\Delta=7$
Next, suppose that $v$ has no support vertex. This means that each $4$-neighbour of $v$ is incident to a $4^{+}$-face containing $v$. Obviously, $v$ has one of the two configurations depicted in Figure 4, where $v_{i},v_{j},v_{k},v_{\ell}$ are $4$-vertices. By Lemma 2.19(a), a $4$-vertex has no 5(5)555(5)5 ( 5 )-neighbours, so the configuration in Figure 4(a) is not possible for $v$. So the neighbours of $v$ can only form as depicted in Figure 4(b). We then infer that $v$ cannot have any $5(4)$-neighbour by Lemma 2.23(b). Thus $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-4\times\frac{1}{4}-\frac{2}{9}-2\times\frac{1}{18}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{2}{9}$ to its 5(5)555(5)5 ( 5 )-neighbour by R7(c), at most $\frac{1}{18}$ to each of its other neighbours by R7(e).
$v$$v_{j}$$v_{k}$$v_{\ell}$$v_{i}$
$v$$v_{j}$$v_{k}$$v_{\ell}$$v_{i}$
@@
(5.2.4). Let $n_{4}(v)=5$. Clearly, $v$ has a support vertex, and so we have $d_{2}(v)\geq 21$ by Lemma 2.26(a). This particularly implies that $v$ has no 5(5)555(5)5 ( 5 )-neighbours by Lemma 2.19(a). Also, $v$ has at most one $5(4)$-neighbour since $d_{2}(v)\geq 21$. Suppose first that $v$ has a $5(4)$-neighbour or two $6(5)$-neighbours. Then we conclude that $v$ is incident to two $5^{+}$-face since $d_{2}(v)\geq 21$. By applying R2, $v$ receives $\frac{1}{5}$ from each of its incident $5^{+}$-faces. In such a case, we have $\mu^{*}(v)\geq 3+2\times\frac{1}{5}-5\times\frac{1}{3}-5\times\frac{1}{4}-\frac{1}{9}-\frac{1}{18}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{1}{9}$ to its $5(4)$-neighbour by R7(d), $\frac{1}{18}$ to each of its $6(5)$-neighbours by R7(e). Now we suppose that $v$ has no $5(4)$-neighbour, and $v$ has at most one $6(5)$-neighbour. Then $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-5\times\frac{1}{4}-\frac{1}{18}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{1}{18}$ to its $6(5)$-neighbour by R7(e).
@@
(5.2.5). Let $n_{4}(v)=6$. Clearly, $v$ has a support vertex, and so we have $d_{2}(v)\geq 21$ by Lemma 2.26(a), which implies that $v$ is incident to two $5^{+}$-faces. In addition, $v$ has no $5(4)$- and 5(5)555(5)5 ( 5 )-neighbours since $d_{2}(v)\geq 21$. By R2, $v$ receives $\frac{1}{5}$ from each of its incident $5^{+}$-faces. Thus we have $\mu^{*}(v)\geq 3+2\times\frac{1}{5}-5\times\frac{1}{3}-6\times\frac{1}{4}-\frac{1}{18}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{1}{18}$ to its $6(5)$-neighbour by R7(e).
@@
(5.3). Let $m_{3}(v)=6$. By Lemma 2.26(d), $v$ has at most five poor neighbours. If $v$ has at most four poor neighbours, then $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-4\times\frac{1}{6}-3\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{6}$ to each of its poor neighbours by R8(a),(c), $\frac{1}{9}$ to each of its $5(4)$-neighbours by R8(b). Suppose further that $v$ has exactly five poor neighbours. First, observe that $v$ has no $4(0)$-neighbours as $m_{3}(v)=6$, i.e., all $4$-neighbours of $v$ are bad $4$-vertices. Note that a $5(4)$-neighbour of $v$ is not adjacent to two poor vertices by Lemma 2.23(a),(b),(c). Moreover, a $4$-vertex has no 5(5)555(5)5 ( 5 )-neighbours by Lemma 2.19(a). Furthermore, a bad $4$-vertex has at most one $4$-neighbour by Lemma 2.19(b). In addition, by Lemma 2.24(b), a 5(5)555(5)5 ( 5 )-vertex is adjacent to at most one 5(5)555(5)5 ( 5 )-vertex. All those facts imply that $v$ has at most one $5(4)$-neighbour. Hence, $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-5\times\frac{1}{6}-\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{6}$ to each of its poor neighbours by R8(a),(c), $\frac{1}{9}$ to its $5(4)$-neighbour by R8(b).
@@
(5.4). Let $m_{3}(v)=7$. By Lemma 2.24(b), a 5(5)555(5)5 ( 5 )-vertex is adjacent to at most one 5(5)555(5)5 ( 5 )-vertex. So, we deduce that $v$ has at most four 5(5)555(5)5 ( 5 )-neighbours. In particular, if $v$ has four such neighbours, then $v$ cannot have any $5(4)$-neighbours by Lemmas 2.23(a) and 2.24(b). Thus $\mu^{*}(v)\geq 3-7\times\frac{1}{3}-4\times\frac{1}{6}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1 and $\frac{1}{6}$ to each of its 5(5)555(5)5 ( 5 )-neighbours by R9(a). If $v$ has three 5(5)555(5)5 ( 5 )-neighbours, then $v$ has at most two $5(4)$-neighbours, and so $\mu^{*}(v)\geq 3-7\times\frac{1}{3}-3\times\frac{1}{6}-2\times\frac{1}{12}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-face by R1, $\frac{1}{6}$ to each of its 5(5)555(5)5 ( 5 )-neighbour by R9(a), and $\frac{1}{12}$ to each of its $5(4)$-neighbours by R9(b). Suppose now that $v$ has at most two 5(5)555(5)5 ( 5 )-neighbours. By Lemma 2.24(b), a 5(5)555(5)5 ( 5 )-vertex has at most one bad $5$-neighbour. We then deduce that $v$ has at most six bad $5$-neighbours when $v$ has at least one 5(5)555(5)5 ( 5 )-neighbour. Thus, $\mu^{*}(v)\geq 3-7\times\frac{1}{3}-2\times\frac{1}{6}-4\times\frac{1}{12}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-face by R1, $\frac{1}{6}$ to each of its 5(5)555(5)5 ( 5 )-neighbour by R9(a), and $\frac{1}{12}$ to each of its $5(4)$-neighbours by R9(b).
@@ 2.3. The case $\Delta=8$
Recall that $G$ does not admit any $2$-distance $23$-coloring, but any planar graph $G^{\prime}$ obtained from $G$ with smaller $|V(G^{\prime})|+|E(G^{\prime})|$ admits a 2-distance $23$-coloring.
Starting with some structural properties of $G$, we address similar results as in the case $\Delta=7$. The proof of the following lemma is omitted since it is similar to Lemma 2.1.
@@
Let $v$ be a $3$-vertex. If $v$ is incident to a $3$-face, say $f_{1}=v_{1}vv_{2}$, then we set $G^{\prime}=G-v+\{v_{1}v_{3}\}$. If $v$ is incident to two $4$-faces, say $f_{1}=v_{1}vv_{2}x$ and $f_{2}=v_{2}vv_{3}y$, then we set $G^{\prime}=G-v+\{v_{1}v_{3}\}$. If $v$ is adjacent to a $6^{-}$-vertex, say $v_{1}$, then we set $G^{\prime}=G-v+\{v_{1}v_{2},v_{1}v_{3}\}$. In each case, $G^{\prime}$ remains proper with respect to $G$ as well as $d_{2}(v)\leq 22$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring with $23$ colors. Since $d_{2}(v)\leq 22$, we can color $v$ with an available color, a contradiction. This particularly implies that $v$ cannot have two (resp. one) $7$-neighbours when $m_{4}(v)=0$ (resp. $m_{4}(v)=1$). ∎
Recall that a $3$-vertex cannot have any $6^{-}$-neighbour by Lemma 2.28(a). That is, a vertex of degree $4$, $5$, or $6$ has no $3$-neighbour. This fact will be assumed in the sequel.
@@
    If $m_{3}(v)=1$, then $v$ has no 5(5)555(5)5 ( 5 )-neighbour. In particular, $v$ has at most one $4$-neighbour.
@@
    If $m_{3}(v)=2$, then $v$ has at most one neighbour consisting of $4$- or $5$-vertices.
@@
Let $v$ be a $4$-vertex.
$(a)$ Let $m_{3}(v)=1$. Obviously, $v$ cannot have any 5(5)555(5)5 ( 5 )-neighbour as $m_{3}(v)=1$. Now, we assume that $v$ has two $4$-neighbours, and so $d_{2}(v)\leq 22$. If we set $G^{\prime}=G-v+\{v_{1}v_{2},v_{1}v_{3},v_{1}v_{4}\}$ for a $4$-neighbour $v_{1}$ of $v$, then $G^{\prime}$ is proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring with $23$ colors. Since $d_{2}(v)\leq 22$, we can color $v$ with an available color, a contradiction.
$(b)$ Let $m_{3}(v)=2$. Suppose that $v$ has two neighbours consisting of $4$- or $5$-vertices. So, $d_{2}(v)\leq 22$, and we get a contradiction as above.
$(c)$ Let $m_{3}(v)\geq 3$. If $v$ has a $4$- or $5(4^{+})$-neighbour, then we have $d_{2}(v)\leq 22$. In a similar way as above, we get a contradiction. ∎
@@
Let $v$ be a $4(1)$-vertex. As $m_{3}(v)=1$, the vertex $v$ is incident to a $3$-face $f_{1}$, say $f_{1}=v_{1}vv_{2}$. Suppose that $m_{4}(v)\geq 2$. By contradiction, assume that $v$ has at most one $7^{+}$-neighbour. This implies that at least one of $v_{1},v_{2}$ is a $6^{-}$-vertex, say $v_{1}$. In particular, we have $d_{2}(v)\leq 22$. If we set $G^{\prime}=G-v+\{v_{1}v_{3},v_{1}v_{4}\}$, then $G^{\prime}$ is proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring with $23$ colors. Since $d_{2}(v)\leq 22$, we can color $v$ with an available color, a contradiction. ∎
@@
    If $m_{4}(v)=0$, then $v$ has two $7^{+}$-neighbours different from 7(7)777(7)7 ( 7 )- and 8(8)888(8)8 ( 8 )-vertices.
@@
    If $m_{4}(v)=1$, then $v$ has either two $8(7^{-})$- or three $7^{+}$-neighbours different from 7(7)777(7)7 ( 7 )- and 8(8)888(8)8 ( 8 )-vertices.
@@
    If $m_{4}(v)=2$, then $v$ has either one $8(7^{-})$- and three $7(6^{-})$-neighbours or two $8(7^{-})$- and one $7(6^{-})$-neighbours or three $8(7^{-})$-neighbours.
@@
Let $v$ be a $4(2)$-vertex, and denote by $f_{1},f_{2},\ldots,f_{4}$ the faces incident to $v$ such that $f_{i}=v_{i}vv_{i+1}$ and $f_{j}=v_{j}vv_{j+1}$ with $i<j$ are $3$-faces. Notice first that if $v$ has a 7(7)777(7)7 ( 7 )- or a 8(8)888(8)8 ( 8 )-neighbours, then there exist two edges in $G[N(v)]$ such that each of them is contained in two $3$-faces.
$(a)$ Let $m_{4}(v)=0$. Assume for a contradiction that $v$ has at most one $7^{+}$-neighbour different from 7(7)777(7)7 ( 7 )- and 8(8)888(8)8 ( 8 )-vertices, i.e., $v$ has three neighbours consisting of $6^{-}$-, 7(7)777(7)7 ( 7 )- or 8(8)888(8)8 ( 8 )-vertices. In such a case, we clearly have $d_{2}(v)\leq 22$. If we set $G^{\prime}=G-v+\{v_{2}v_{4}\}$ (when $i=1$, $j=2$) or $G^{\prime}=G-v+\{v_{2}v_{3},v_{1}v_{4}\}$ (when $i=1$, $j=3$), then $G^{\prime}$ is proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring with $23$ colors. Since $d_{2}(v)\leq 22$, we can color $v$ with an available color, a contradiction.
$(b)$ Let $m_{4}(v)=1$. Notice that $v$ cannot have two neighbours consisting of 7(7)777(7)7 ( 7 )- or 8(8)888(8)8 ( 8 )-vertices as $m_{3}(v)=2$. Suppose first that $v$ has no 7(7)777(7)7 ( 7 )- and 8(8)888(8)8 ( 8 )-vertices. If $v$ has either two $6^{-}$- and one $7$-neighbours or three $6^{-}$-neighbours, then we would have $d_{2}(v)\leq 22$, and similarly as above, we get a contradiction. Thus, we conclude that $v$ has either two $8(7^{-})$-neighbours or three $7^{+}$-neighbours. Now we suppose that $v$ has a 7(7)777(7)7 ( 7 )- or 8(8)888(8)8 ( 8 )-neighbours. If $v$ has two $6^{-}$-neighbours, then we have $d_{2}(v)\leq 21$. Moreover, if $v$ has both a $6^{-}$-neighbour and a $7$-neighbour different from 7(7)777(7)7 ( 7 )-vertex, then we have $d_{2}(v)\leq 22$. Similarly as above, we get a contradiction. Thus, the claim holds.
$(c)$ Let $m_{4}(v)=2$. We first show that $v$ has at least one $8(7^{-})$-neighbours. Indeed, if all neighbours of $v$ are $7^{-}$- or 8(8)888(8)8 ( 8 )-vertices, then we have $d_{2}(v)\leq 22$ since $m_{4}(v)=2$. In a similar way as above, we get a contradiction. Therefore, $v$ has at least one $8(7^{-})$-neighbours. If $v$ has exactly one $8(7^{-})$-neighbours, and exactly two $7(6^{-})$-neighbours, or if $v$ has exactly two $8(7^{-})$-neighbours, and no $7(6^{-})$-neighbours, then we have $d_{2}(v)\leq 22$. Similarly as above, we get again a contradiction. ∎
@@
    If $m_{4}(v)=0$, then $v$ has either one $8(7^{-})$- and three $7(6^{-})$-neighbours or two $8(7^{-})$- and one $7(6^{-})$-neighbours or two $8(7^{-})$- and two 8(8)888(8)8 ( 8 )-neighbours or three $8(7^{-})$-neighbours.
@@
    If $m_{4}(v)=1$, then $v$ has either two $8(7^{-})$- and two $7(6^{-})$-neighbours or three $8(7^{-})$-neighbours.

### v3

@@ 2.2. The case $\Delta=7$
Next, suppose that $v$ has no support neighbour. This means that each $4$-neighbour of $v$ is incident to a $4^{+}$-face containing $v$. Obviously, $v$ has one of the two configurations depicted in Figure 4, where $v_{i},v_{j},v_{k},v_{\ell}$ are $4$-vertices. By Lemma 2.21(a), a $4$-vertex has no $5(5)$-neighbours, so the configuration in Figure 4(a) is not possible for $v$. So the neighbours of $v$ can only form as depicted in Figure 4(b). We then infer that $v$ cannot have any $5(4)$-neighbour by Lemma 2.25(b). Thus $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-4\times\frac{1}{4}-\frac{2}{9}-2\times\frac{1}{18}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{2}{9}$ to its $5(5)$-neighbour by R7(c), at most $\frac{1}{18}$ to each of its other neighbours by R7(e).
[…]
@@
(5.2.4). Let $n_{4}(v)=5$. Then $v$ has no $5(5)$-neighbours by Lemma 2.21(a). Also, $v$ has at most one $5(4)$-neighbour by Lemma 2.28(i). Suppose first that $v$ has a $5(4)$-neighbour or two $6(5)$-neighbours. Then $v$ is incident to two $5^{+}$-face by Lemma 2.28(i). By applying R2, $v$ receives $\frac{1}{5}$ from each of its incident $5^{+}$-faces. In such a case, we have $\mu^{*}(v)\geq 3+2\times\frac{1}{5}-5\times\frac{1}{3}-5\times\frac{1}{4}-\frac{1}{9}-\frac{1}{18}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{1}{9}$ to its $5(4)$-neighbour by R7(d), $\frac{1}{18}$ to each of its $6(5)$-neighbours by R7(e). Now we suppose that $v$ has no $5(4)$-neighbour, and $v$ has at most one $6(5)$-neighbour. Then $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-5\times\frac{1}{4}-\frac{1}{18}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{1}{18}$ to its $6(5)$-neighbour by R7(e).
@@
(5.2.5). Let $n_{4}(v)=6$. By Lemma 2.28(j), $v$ is incident to two $5^{+}$-faces, in particular, $v$ has no $5(4)$- or $5(5)$-neighbours. By R2, $v$ receives $\frac{1}{5}$ from each of its incident $5^{+}$-faces. Thus we have $\mu^{*}(v)\geq 3+2\times\frac{1}{5}-5\times\frac{1}{3}-6\times\frac{1}{4}-\frac{1}{18}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{1}{18}$ to its $6(5)$-neighbour by R7(e).
@@
(5.3). Let $m_{3}(v)=6$. By Lemma 2.28(d), $v$ has at most five poor neighbours. If $v$ has at most four poor neighbours, then $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-4\times\frac{1}{6}-3\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{6}$ to each of its poor neighbours by R8(a),(c), $\frac{1}{9}$ to each of its $5(4)$-neighbours by R8(b). Suppose further that $v$ has exactly five poor neighbours. First, observe that $v$ has no $4(0)$-neighbours as $m_{3}(v)=6$, i.e., all $4$-neighbours of $v$ are bad $4$-vertices. Note that, by Lemma 2.25(a),(b), a $5(4)$-neighbour of $v$ is not adjacent to two $5(5)$-vertices, and is not adjacent to both a $5(5)$- and $4$-vertices. Moreover, a $4$-vertex has no $5(5)$-neighbours by Lemma 2.21(a). Furthermore, a bad $4$-vertex has at most one $4$-neighbour by Lemma 2.21(b), and has no both $4$- and $5(4)$-neighbour by Lemma 2.21(c). In addition, by Lemma 2.26(b), a $5(5)$-vertex is adjacent to at most one $5(5)$-vertex. All those facts imply that $v$ has at most one $5(4)$-neighbour. Hence, $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-5\times\frac{1}{6}-\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{6}$ to each of its poor neighbours by R8(a),(c), $\frac{1}{9}$ to its $5(4)$-neighbour by R8(b).
@@
(5.4). Let $m_{3}(v)=7$. By Lemma 2.26(b), a $5(5)$-vertex is adjacent to at most one $5(5)$-vertex. So, we deduce that $v$ has at most four $5(5)$-neighbours. In particular, if $v$ has four such neighbours, then $v$ cannot have any $5(4)$-neighbours by Lemmas 2.25(a) and 2.26(b). Thus $\mu^{*}(v)\geq 3-7\times\frac{1}{3}-4\times\frac{1}{6}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1 and $\frac{1}{6}$ to each of its $5(5)$-neighbours by R9(a). If $v$ has three $5(5)$-neighbours, then $v$ has at most two $5(4)$-neighbours by Lemmas 2.25(a) and 2.26(b), and so $\mu^{*}(v)\geq 3-7\times\frac{1}{3}-3\times\frac{1}{6}-2\times\frac{1}{12}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-face by R1, $\frac{1}{6}$ to each of its $5(5)$-neighbour by R9(a), and $\frac{1}{12}$ to each of its $5(4)$-neighbours by R9(b). Suppose now that $v$ has at most two $5(5)$-neighbours. By Lemma 2.26(b), a $5(5)$-vertex has at most one bad $5$-neighbour. We then deduce that $v$ has either seven $5(4)$-neighbours or at most six bad $5$-neighbours. In the worst case, $\mu^{*}(v)\geq 3-7\times\frac{1}{3}-2\times\frac{1}{6}-4\times\frac{1}{12}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-face by R1, $\frac{1}{6}$ to each of its $5(5)$-neighbour by R9(a), and $\frac{1}{12}$ to each of its $5(4)$-neighbours by R9(b).
@@ 2.3. The case $\Delta=8$
Recall that $G$ does not admit any $2$-distance $23$-coloring, whereas any planar graph $G^{\prime}$ obtained from $G$ with smaller value of $|V(G^{\prime})|+|E(G^{\prime})|$ admits a $2$-distance $23$-coloring.
We begin by establishing several structural properties of $G$, analogous to those obtained for the cases $\Delta=6$ and $\Delta=7$. The proof of the following lemma is omitted, as it follows from the same arguments used in Lemma 2.3.
@@
If $v$ is incident to a $3$-face, say $f_{1}=v_{1}vv_{2}$, then we set $G^{\prime}=G-v+\{v_{1}v_{3}\}$. If $v$ is incident to two $4$-faces, say $f_{1}=v_{1}vv_{2}x$ and $f_{2}=v_{2}vv_{3}y$, then we again set $G^{\prime}=G-v+\{v_{1}v_{3}\}$. If $v$ is adjacent to a $7^{-}$-vertex, say $v_{1}$, then we set $G^{\prime}=G-v+\{v_{1}v_{2},v_{1}v_{3}\}$. In each case, $G^{\prime}$ remains proper with respect to $G$, and moreover $d_{2}(v)\leq 22$ by Lemma 2.2. This yields a contradiction by Remark 2.1. Thus, none of the above configurations can occur. ∎
Since a $3$-vertex cannot have any $6^{-}$-neighbour by Lemma 2.30(a), it follows that every vertex of degree $4$, 55, or $6$ has no $3$-neighbour. This fact will be assumed in the remainder of this section.
@@
    If $m_{3}(v)=1$, then $v$ has no $5(5)$-neighbour. In particular, $v$ has at most one $4$-neighbour.
@@
    If $m_{3}(v)=2$, then $v$ has at most one $5^{-}$-neighbour.
@@
$(a)$ Let $m_{3}(v)=1$. Clearly, $v$ cannot have any $5(5)$-neighbour since $m_{3}(v)=1$. Now assume that $v$ has two $4$-neighbours. Then $d_{2}(v)\leq 22$ by Lemma 2.2. If we set $G^{\prime}=G-v+\{v_{1}v_{2},v_{1}v_{3},v_{1}v_{4}\}$ for a $4$-neighbour $v_{1}$ of $v$, then $G^{\prime}$ is proper with respect to $G$. By Remark 2.1, this yields a contradiction.
$(b)$ Let $m_{3}(v)=2$. Suppose that $v$ has two neighbours that are $4$- or $5$-vertices. Then $d_{2}(v)\leq 22$ by Lemma 2.2, and the same reduction as in $(a)$ leads to a contradiction.
$(c)$ Let $m_{3}(v)\geq 3$. If $v$ has a $4$-neighbour or a $5(4^{+})$-neighbour, then again $d_{2}(v)\leq 22$ by Lemma 2.2. As in the previous cases, this contradicts Remark 2.1. ∎
@@
Since $m_{3}(v)=1$, the vertex $v$ is incident to a $3$-face $f_{1}=v_{1}vv_{2}$. Assume that $m_{4}(v)\geq 2$. Suppose, for a contradiction, that $v$ has at most one $7^{+}$-neighbour. Then at least one of $v_{1},v_{2}$ must be a $6^{-}$-vertex, say $v_{1}$. In particular, $d_{2}(v)\leq 22$ by Lemma 2.2. If we set $G^{\prime}=G-v+\{v_{1}v_{3},v_{1}v_{4}\}$, then $G^{\prime}$ is proper with respect to $G$, a contradiction to Remark 2.1. ∎
@@
    If $m_{4}(v)=0$, then $v$ has two $7^{+}$-neighbours that are neither $7(7)$- nor $8(8)$-vertices.
@@
    If $m_{4}(v)=1$, then $v$ has either two $8(7^{-})$-neighbours or three $7^{+}$-neighbours different from $7(7)$- and $8(8)$-vertices.
@@
    If $m_{4}(v)=2$, then $v$ has one of the following: one $8(7^{-})$- and three $7(6^{-})$-neighbours, or two $8(7^{-})$- and one $7(6^{-})$-neighbour, or three $8(7^{-})$-neighbours.
@@
Let $f_{i}=v_{i}vv_{i+1}$ and $f_{j}=v_{j}vv_{j+1}$ with $i<j$ be the two $3$-faces incident to $v$. Notice first that if $d_{2}(v)\leq 22$, then we set $G^{\prime}=G-v+\{v_{2}v_{4}\}$ when $(i,j)=(1,2)$; we set $G^{\prime}=G-v+\{v_{2}v_{3},\,v_{1}v_{4}\}$ when $(i,j)=(1,3)$. In each case, $G^{\prime}$ is proper with respect to $G$. By Remark 2.1, this yields a contradiction. Thus we further assume that $d_{2}(v)\geq 23$. We also note that if $v$ has a $7(7)$- or $8(8)$-neighbour, then at least two edges in $E(v)$ lie in two $3$-faces. Moreover, $v$ cannot have two neighbours consisting of $7(7)$-, or $8(8)$-vertices since $m_{3}(v)=2$.
$(a)$ Let $m_{4}(v)=0$. Assume for a contradiction that $v$ has at most one $7^{+}$-neighbour that is not $7(7)$- or $8(8)$-vertices; equivalently, $v$ has three neighbours consisting of $6^{-}$-, $7(7)$-, or $8(8)$-vertices. Then $d_{2}(v)\leq 22$ by Lemma 2.2, a contradiction.
$(b)$ Let $m_{4}(v)=1$. First suppose that $v$ has no $7(7)$- or $8(8)$-neighbours. If $v$ has either two $6^{-}$-neighbours and one $7$-neighbour, or three $6^{-}$-neighbours, then $d_{2}(v)\leq 22$ by Lemma 2.2, giving a contradiction. Hence $v$ has either two $8(7^{-})$-neighbours or three $7^{+}$-neighbours. Now suppose $v$ has a $7(7)$- or $8(8)$-neighbour. If $v$ has two $6^{-}$-neighbours, then $d_{2}(v)\leq 21$, and if $v$ has one $6^{-}$- and one $7$-neighbour (other than $7(7)$), then $d_{2}(v)\leq 22$; in both cases we reach a contradiction as before. Thus the statement follows.
$(c)$ Let $m_{4}(v)=2$. We first show that $v$ has at least one $8(7^{-})$-neighbour. Indeed, if all neighbours of $v$ were $7^{-}$- or $8(8)$-vertices, then $d_{2}(v)\leq 22$ by Lemma 2.2, a contradiction. Thus $v$ has at least one $8(7^{-})$-neighbour. If $v$ has exactly one $8(7^{-})$-neighbour and exactly two $7(6^{-})$-neighbours, then $d_{2}(v)\leq 22$. Likewise, if $v$ has exactly two $8(7^{-})$-neighbours and no $7(6^{-})$-neighbours, then again $d_{2}(v)\leq 22$. In each case we obtain a contradiction. Therefore, the remaining possibilities are those listed in the statement: $v$ has either one $8(7^{-})$- and three $7(6^{-})$-neighbours, or two $8(7^{-})$- and one $7(6^{-})$-neighbour, or three $8(7^{-})$-neighbours. ∎
@@
    If $m_{4}(v)=0$, then $v$ has one of the following: one $8(7^{-})$- and three $7(6^{-})$-neighbours, or two $8(7^{-})$- and one $7(6^{-})$-neighbour, or two $8(7^{-})$- and two $8(8)$-neighbours, or three $8(7^{-})$-neighbours.
@@
    If $m_{4}(v)=1$, then $v$ has either two $8(7^{-})$- and two $7(6^{-})$-neighbours, or three $8(7^{-})$-neighbours.